# A Tale of Single-channel Electroencephalogram: Devices, Datasets, Signal Processing, Applications, and Future Directions

### v1

@@ VI-E Other Applications
The single-channel EEG with steady state visually evoked potential (SSVEP) paradigm demonstrates significant potential in BCI, neuroscience research and clinical diagnosis. Single-channel EEG devices typically record signals in the occipital region, as it is most sensitive to visual stimuli. Nguyen et al. developed a SSVEP speller system based on monopolar single-channel EEG, using a one-dimensional CNN to detect SSVEP frequency and a custom head-mounted device (O1-Oz channel) with a virtual keyboard containing 58 characters. In experiments with eight subjects, the system achieved an average classification accuracy of 99.2%, and in actual online spelling experiments, it achieved an accuracy of 97.4% and an information transmission rate of 49 ± 7.7 bits/min [155]. Karunasena et al. proposed a streamlined SSVEP-based BCI system that utilizes a single-channel EEG to control a robotic arm. By capturing the brain’s visual response to LEDs flashing at different frequencies, the system generates control commands by focusing on LEDs of specific frequencies, effectively controlling the robotic arm [156]. The single-channel EEG with motor imagery paradigm is also a significant area of BCI research. Ge et al. used single-channel EEG to classify four types of motor imagery, decomposing the signal into the time-frequency domain through STFT, constructing multi-channel information, and combining the common spatial pattern algorithm and SVM to achieve high classification accuracy from the sensorimotor and frontal areas [157]. Due to its user-friendliness, convenience, and low cost, single-channel EEG has also been applied in microsleep detection to enhance the safety of intelligent transportation systems (ITS). Chougule et al. developed a deep learning model named Diamond Net, which uses an attention mechanism combined with WT and STFT spectrograms to effectively distinguish microsleep from other sleep stages [158].
@@ VII DISCUSSION
Single-channel EEG is pivotal due to its low cost, comfort, and non-invasiveness. The increasing number of research articles highlights its growing significance, reflecting technological advancements and a shift toward accessible brain monitoring solutions. Researchers, consumers, and clinicians are increasingly adopting single-channel EEG in various settings, such as neuroscience exploration, human-computer interaction, education, mental health monitoring, and clinical diagnosis. Firstly, a critical aspect of single-channel EEG’s utility lies in its signal processing and data handling capabilities. Traditionally, EEG involved multiple channels for comprehensive brain activity mapping, but single-channel systems require sophisticated computational techniques to enhance accuracy and interpretability. Secondly, the evolution and innovation of single-channel EEG devices in portability, reduced setup complexity, rapid deployment, and continuous monitoring are crucial for widespread adoption. Technological advancements in hardware and software, including miniaturization, sensor sensitivity, and wireless capabilities, are expected to enhance signal quality and reliability. Sophisticated ML algorithms are essential for robust signal processing and data analysis in single-channel EEG systems. Thirdly, significant challenges such as ensuring high signal quality and overcoming the inherent limitations of single-channel systems, like data depth and interpretative consistency, are expected to be further addressed through AI-based generative models [159]. AI-based EEG generation techniques have demonstrated the potential to generate multi-channel signals with high quality from single-channel signals, matching or even surpassing the performance of multichannel systems. Finally, rigorous validation and standardization of generative AI models in single-channel EEG systems are crucial, along with developing robust frameworks to navigate ethical and privacy concerns.
@@ VIII CONCLUSION
Single-channel EEG, benefiting from its low cost, comfort, and non-invasiveness, holds great potential for real-time brain activity monitoring. The increasing number of research articles underscores its growing significance. This review comprehensively covers the development trends, devices, datasets, signal processing methods, applications, and future directions of single-channel EEG, serving as a guide for influential and novel research to further advance single-channel EEG systems. Moreover, this review clarifies the definitions of unipolar and bipolar configurations in single-channel EEG. Wired single-channel EEG systems capture significant information with minimal channels, facilitating tasks such as sleep staging and disease detection. Wireless single-channel EEG systems, noted for their user-friendliness and quick setup, are widely utilized in education research, emotion recognition, and other fields. The unique advantages of single-channel EEG enable its success in diverse application environments, and will continue to play a pivotal role in driving the development and innovation of neuroscience technology.

### v2

@@ VI-F Other Applications
The single-channel EEG with SSVEP paradigm demonstrates significant potential in BCI, neuroscience research and clinical diagnosis, with devices typically recording signals in the occipital region for visual stimuli sensitivity. Nguyen et al. developed a SSVEP speller system using monopolar single-channel EEG with one-dimensional CNN and a custom head-mounted device, achieving 99.2% classification accuracy in experiments and 97.4% accuracy with 49 ± 7.7 bits/min transmission rate in online spelling tests [184]. Karunasena et al. demonstrated a streamlined SSVEP-based BCI system controlling a robotic arm through visual responses to frequency-specific LED stimuli [185]. The single-channel EEG with motor imagery paradigm is also a significant area of BCI research. Ge et al. used single-channel EEG to classify four types of motor imagery, decomposing the signal into the time-frequency domain through STFT, constructing multi-channel information, and combining the common spatial pattern algorithm and SVM to achieve high classification accuracy from the sensorimotor and frontal areas [186]. Due to its user-friendliness, convenience, and low cost, single-channel EEG has also been applied in microsleep detection to enhance the safety of intelligent transportation systems (ITS). Chougule et al. developed Diamond Net, which uses an attention mechanism combined with WT and STFT spectrograms to effectively distinguish microsleep from other sleep stages [187].
Figure 6: Advancements in single-channel EEG hardware focus on enhancing sensitivity, integration, comfort, and energy efficiency to improve device performance and user experience.
@@ VII DISCUSSION
Single-channel EEG demonstrates pivotal advantages through low cost, comfort, and non-invasiveness. The proliferation of research articles reflects its growing significance, driven by technological advancements toward accessible brain monitoring. Its adoption spans diverse applications from neuroscience exploration to clinical diagnosis, including human-computer interaction, education, and mental health monitoring. Firstly, a critical aspect of single-channel EEG’s utility lies in its signal processing and data handling capabilities. Traditionally, EEG involved multiple channels for comprehensive brain activity mapping, but single-channel systems require sophisticated computational techniques to enhance accuracy and interpretability. Secondly, single-channel EEG innovation emphasizes portability, reduced setup complexity, rapid deployment, and continuous monitoring capabilities. Advances in hardware and software through miniaturization, enhanced sensor sensitivity, and wireless capabilities promise improved signal quality and reliability, while sophisticated ML algorithms enable robust signal processing and data analysis. Thirdly, significant challenges such as ensuring high signal quality and overcoming the inherent limitations of single-channel systems, like data depth and interpretative consistency, are expected to be further addressed through AI-based generative models. AI-based EEG generation techniques have demonstrated the potential to generate multi-channel signals with high quality from sparse-channel signals, matching or even surpassing the performance of multichannel systems [188]. Finally, rigorous validation and standardization of generative AI models in single-channel EEG systems are crucial, along with developing robust frameworks to navigate ethical and privacy concerns. Several critical areas of development are anticipated to influence the future trajectory of single-channel EEG technology.
VII-A Advanced Signal Processing with AI Integration
Recent advancements in signal processing and AI have transformed single-channel EEG applications through technological innovations: CNNs demonstrate exceptional spatial feature extraction capabilities [189], one-dimensional CNNs excel in mental state classification and anomaly detection [190], while LSTM networks effectively model temporal dependencies [191], as evidenced by correlation-based epileptic signal classification methodology [192]. The field’s progressive evolution encompasses sophisticated transfer learning strategies enabling cross-domain knowledge application in sleep staging scenarios [193], advanced generative models including Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) facilitating data augmentation and reconstruction [194, 195], innovative auxiliary synthesis frameworks bridging the gap between single and multi-channel systems [196], refined adaptive filtering techniques for artifact mitigation in dynamic environments [197, 198], and seamless edge computing integration enabling continuous monitoring and real-time neurofeedback applications [199].
Future advancements in single-channel EEG technology will be fundamentally driven by specialized DL architectures optimized for single-channel data characteristics, enhancing both classification accuracy and computational efficiency. The evolution of transfer learning methodologies will enable effective adaptation of multi-channel-trained models to single-channel applications, incorporating advanced domain adaptation techniques [200]. The integration of generative models will extend beyond multi-channel data reconstruction to enhance signal quality and generate synthetic training datasets. Concurrent development of real-time signal processing algorithms will facilitate sophisticated artifact rejection and noise reduction mechanisms for wearable devices, ensuring robust EEG measurements in dynamic environments [35]. Moreover, advances in edge AI and on-device learning will empower wearable EEG systems with embedded intelligence, minimizing latency while preserving data privacy through reduced external transmission requirements [201, 202, 203]. Collectively, these synergistic advancements will not only elevate the performance and applicability of single-channel EEG systems but also enhance their accessibility, thereby fostering broader adoption across clinical, consumer, and research domains.
VII-B Technological Innovations in Hardware
Figure 7: Integrating single-channel EEG data with wearable technology and IoT enables early disease prediction and alerts, facilitating timely intervention by sharing insights with physicians and family members.
Hardware advancements have been fundamental to the evolution of single-channel EEG devices, with innovations primarily centered on four critical aspects: sensitivity, integration, comfort, and energy efficiency, as illustrated in Fig. 6. A typical EEG recording system comprises an amplifier, low pass filter, and ADC [204], where the amplifier serves as the crucial interface between electrode and signal processing engine. Contemporary high-gain amplifiers utilize low-noise architectures to achieve signal amplification while minimizing noise introduction. Marzieh et al. developed an innovative two-stage amplifier incorporating composite transistors and feedback structure, demonstrating superior direct current gain and slew rate, while achieving lower input reference noise compared to conventional metal-oxide-semiconductor (CMOS) amplifiers [205]. Furthermore, Cheng et al. empirically validated that strategic selection of compensation capacitors significantly enhances amplifier performance, facilitating more precise brainwave signal acquisition [206]. Beyond amplifier innovations, electrode design has evolved significantly. Novel semi-dry and dry contact electrodes have emerged as promising alternatives to conventional Ag/AgCl wet electrodes [207]. Semi-dry electrodes combine advantages of wet and dry variants, delivering superior signal recording with minimal electrolyte usage, though their application remains limited to short-term monitoring due to electrolyte drying [208]. Dry electrodes excel in long-term monitoring while offering enhanced user comfort, particularly suitable for wireless wearable single-channel EEG devices. A critical frontier in electrode advancement lies in developing sophisticated materials to minimize contact impedance, with metals, carbon-based compounds, oxides, and polymers offering diverse fabrication possibilities [209]. Furthermore, the integration of flexible electronic technologies represents a pivotal research direction. While these technologies have predominantly been implemented in wet electrode systems [210], their application in long-term dry electrode monitoring remains an active area of investigation, with numerous studies currently exploring this promising avenue [211].
System on Chip (SoC) design represents a pivotal innovation in enhancing the integration of single-channel EEG devices. By consolidating multiple functional modules, such as signal acquisition, processing, and communication, onto a single chip, manufacturers can significantly reduce the device’s footprint while improving overall performance. Mohammad et al. introduced a low-power, high-input-impedance EEG signal acquisition SoC that integrates an instrumentation amplifier, a neural signal-specific analog-to-digital converter (NSS-ADC), active electrodes, a transconductance-driven right leg, and a common-mode feedback module. This integration demonstrates its effectiveness and advantages in continuously acquiring EEG signals for wearable applications [212]. Additionally, there is a growing interest in the integration of EEG with other modalities, particularly functional near-infrared spectroscopy (fNIRS), for a multimodal assessment of brain function. The EEG-fNIRS combination can provide a comprehensive picture of both electrophysiological and hemodynamic processes with high spatiotemporal resolution. The development of wearable, mechanically, and electrically integrated EEG-fNIRS technologies is a crucial next step in advancing this field [213]. Furthermore, energy-efficient processors represent a significant innovation enabling prolonged EEG data collection and real-time analysis. Research advances in ultra-low-power technologies utilizing human body energy through heat or motion [214] promise to eliminate battery dependency, advancing sustainable EEG monitoring capabilities..
VII-C Integration with Wearable Technology and IoT
The proliferation of wearable devices and Internet of Things (IoT) technology has revolutionized EEG signal acquisition, enabling more streamlined and naturalistic monitoring processes while facilitating intelligent automation through integrated detection systems [215]. Traditional wired EEG systems present significant limitations in terms of user experience and practical applicability [216]. In contrast, recent innovations in wireless EEG systems, particularly in wearable devices, enhance portability and user flexibility, advancing from conventional fixed systems toward intelligent wireless solutions prioritizing comfort and signal fidelity. While wearable devices present data quality and motion artifact challenges, their cost-effectiveness, wireless connectivity, lightweight construction, and operational simplicity, combined with continuous hardware improvements, demonstrate increasing potential and performance. Fig. 7 demonstrates how the integration of single-channel EEG data with wearable technology and IoT infrastructure enables predictive disease detection and early warning systems, facilitating timely medical intervention through seamless information sharing among healthcare providers and family members.
The convergence of wearable devices and IoT technology facilitates specialized applications leveraging EEG monitoring, with detection systems demonstrating IoT device control capabilities [217]. Single-channel EEG systems present a compelling integration trajectory through their minimal form factor and cost-effectiveness. While conventional multi-channel EEG systems deliver comprehensive neurological data, their substantial footprint limits practical deployment. Despite reduced spatial resolution, single-channel EEG devices employ advanced signal processing and algorithms for substantive neurophysiological information extraction, enabling mobile applications in real-time monitoring and personalized IoT environments. Research demonstrates applications in driver fatigue monitoring [218] and IoT-integrated seizure prediction through cloud services [219], implementing DL algorithms for prediction and automated healthcare notifications. IoT infrastructure capabilities [220] with centralized cloud analytics enable efficient detection and rapid emergency response protocols. As technological capabilities evolve, single-channel EEG systems are projected to enhance the accessibility of neurological monitoring within IoT-enabled environments.
@@ VIII CONCLUSION
Single-channel EEG, benefiting from its low cost, comfort, and non-invasiveness, holds great potential for real-time brain activity monitoring. The increasing number of research articles underscores its growing significance. This comprehensive review has analyzed the evolution and landscape of single-channel EEG, revealing advances in hardware design, signal processing methodologies, and applications. Consumer-grade devices have democratized brain-computer interaction, while sophisticated signal processing and ML algorithms have enhanced signal quality and feature extraction. These improvements have enabled robust applications across sleep analysis, emotion recognition, and clinical diagnostics, particularly for neurological conditions such as depression, AD, MCI, and epilepsy. AI integration has further expanded these capabilities, enabling more accurate classification and real-time analysis. The convergence of advanced hardware, signal processing, and AI presents opportunities for innovation, while future developments in AI-enhanced processing, smart hardware, and IoT connectivity will drive the next generation of applications, advancing the accessibility and reliability of brain monitoring technologies. We believe this review of single-channel EEG will facilitate future research and innovation in brain monitoring systems.